# From Agent Simulation to Social Simulator: A Comprehensive Review (Part 2)

### v1

@@ I The Laws of System Complexity
Complexity is one of the most intricate concepts in modern science [7, 8], and no unified definition has yet been established; perhaps such a definition may never exist. The different layers of the real world—physical, biological, social, and conscious layers—each exhibit distinct forms of complexity that should not be conflated. One cannot substitute lower-level complexity for higher-level complexity, nor dismiss lower-level complexity on the basis of higher-level complexity. Complexity can be understood from various perspectives and at different levels. This section first provides a categorical overview of the causes of complexity, followed by a detailed discussion using structural complexity and distributional complexity as examples. The formulas presented here are mainly drawn from Claudio Cioffi-Revilla’s work Introduction to Computational Social Science: Principles and Applications. For more details, please refer to the original text [9].
I-A Emergent Phenomena in Complex Systems
@@
Complexity Arising from Individual Intelligence: Even if a system composed of non-intelligent individuals generates complexity through self-organization, it is usually relatively rudimentary and low-level. Intelligent agent systems (such as those studied by the Santa Fe Institute’s CAS research) can sense their surrounding environments, predict future system states, learn from experience, and develop behavioral rules, thereby achieving adaptive transformations. Thus, they are inherently complex. The higher the intelligence of individual agents, the greater the system’s complexity. Agent intelligence is one of the major sources of complexity. For example, land combat systems typically involve a limited number of agents but are still complex due to the high intelligence of each agent, resulting in unpredictable and emergent system behaviors. These systems are therefore classified as open complex systems.
@@ II-A Horizontal Causation and Vertical Causation
In contrast, vertical causality argues that opening the black box involves decomposing the system into its individual components and generating the macroscopic phenomena under study by reconstructing the dynamic interactions between these components. Based on this perspective, priority should be given to methods that provide credible narratives to explain observed patterns. This view maintains that dependencies do not constitute causal mechanisms themselves, but rather are manifestations of such mechanisms. In response to this viewpoint, simulation methods become powerful tools for investigating the complex dynamic details between events. In particular, Agent-Based Modeling (ABM) is regarded as a key tool for establishing causal mechanisms, as it can demonstrate that the dependencies under study can be derived from hypothetical dynamic narratives [24].
@@
In contrast, the ”vertical causation” camp has embraced theory-driven approaches as its mainstream, requiring macroeconomic models to be integrated with ”microfoundations.” Initially, this led to the widespread adoption of rational choice models, which calculate economic aggregates based on individual preferences and expectations [27, 28]. Subsequently, Agent-Based Modeling (ABM) was introduced to address issues related to agent heterogeneity and the aggregation of social networks [29, 30, 31]. Sociology—analytical sociology in particular—has followed a similar path: shifting from regression-based statistical methods for survey data to ABM, in order to tackle micro-to-macro problems when interdependent structures exist [32].
@@
From the perspective of vertical causation, however, this implication is incorrect, because the role of variables and functions in statistical models differs from their role in Agent-Based Modeling (ABM). In the former case, variables and functions are used to detect patterns of average effects—patterns that may reflect the overall statistical characteristics of the hypothesized underlying mechanisms (composed of entities, attributes, and interactions). In contrast, in the latter case, variables and functions are employed to precisely represent the attributes, activities, and interactions of relevant micro-level entities. Furthermore, functions are used to initiate (i.e., simulate) the dynamics associated with the hypothesized mechanisms, with the goal of assessing under what conditions such dynamics give rise to the relevant macro-level phenomena.
@@ II-B Forward Causation and Reverse Causation
Goldthorpe further clarified that causal relationships can be understood in at least three distinct ways: robust dependence, manipulability of outcomes, and generative processes [38]. In the first view, a causal claim depends on whether X continues to affect Y after introducing a set of other variables (denoted as Z) into the analysis. This perspective sees causality as contingent on controlling for confounding variables. Methods aligned with this view include time-series analysis, early causal path analysis, structural equation models (SEM), and more broadly, regression-based multivariate analytical techniques used in survey data analysis. In the second view, causality is interpreted as follows: if a causal factor X is manipulated, then under appropriate controls of confounding variables, a systematic effect on the response variable Y will be observed. This manipulative view of causality is aligned with the methodology of randomized experiments. Lastly, when causation is understood as a generative process, the focus shifts to explaining the nature and validity of the relational processes on which causality depends. In this view, simulation methods are considered promising tools for explaining underlying mechanisms.
@@
The third scenario belongs to the reverse causal research paradigm, characterized by a longitudinal, generative perspective, and often operationalized through Agent-Based Modeling (ABM). In this framework, Y denotes the social phenomenon to be explained, X refers to a set of agents with diverse heterogeneous attributes, and rules specify agent behaviors and interaction mechanisms. Simulations illustrate how micro-level interactions give rise to specific macro-level phenomena. The actions and interactions of agents are the causes of emergent macro patterns. Tracing how agent-level behaviors and relationships “step-by-step” lead to these patterns provides a mechanistic explanation of the observed outcomes.
@@ III-A Computational Experiment = ABM + Experiment
In response to the challenges faced by ABM in the field of causal inference, Professor Fei-Yue Wang systematically proposed the basic ideas, concepts and methods of computational experiment method in 2004, emphasizing the circular feedback relationship between artificial systems and actual systems [39, 40, 41]. In recent years, Professor Xue and his team have carried out extensive and systematic work on computational experiment methods [42, 43, 44, 45, 46, 47, 48, 49, 50, 48, 51]. They provides a general framework for computational experiment method, which can integrate ABM logic and experimental techniques, and support the customization of general method for different application fields [6, 52, 53]. The framework of computational experiment method includes five steps: artificial society modeling, experimental system construction, computational experiment design, computational experiment analysis, and computational experiment verification, and forms a feedback closed loop [54, 55, 56, 57, 58].
Computational experiment method are an improved and extended version of ABM. ABM is not simply used as a simulation tool, but as an “artificial social laboratory” that “spontaneously grows and cultivates” an alternative version of the actual system, and can conduct various “counterfactual experiments” on system behavior and decision analysis. On the one hand, computational experiments use ABM to emphasize vertical causality [59], pay attention to the causes and formation process of phenomena, and give causal significance to the cross-level correlation information generated by ABM (the connection between micro-mechanisms and macro-emergence). On the other hand, computational experiments “use counterfactual algorithms to construct multiple worlds” to emphasize horizontal causality, establish a probabilistic relationship between model parameters and macro-emergence through experimental mechanisms, reveal the patterns and laws within the system, and provide “reference”, “prediction” and “guidance” for the possible operation of the actual system. Because the virtual foundation on which the experimental mechanism is based, that is, the attributes, behaviors and local environment of the agent, is a replica of its real-world counterpart. Therefore, the probabilistic relationships established between different levels of analysis can be viewed as counterfactual connections across levels of analysis in the real world.
@@ III-A3 Scenario deduction
III-A4 Traceability analysis
@@ III-B2 Explanatory Module: Factorial Experiment Design
III-B3 Predictive Module: Meta-Model of Artificial Society
@@ III-C3 Mechanism Analysis (Linking Micro and Macro)
The third layer represents the reverse causation research paradigm. This level involves not only analyzing the objective world model but also focusing on the agents’ subjective world models. Here, Y denotes the social phenomenon to be explained; X represents a set of agents with heterogeneous attributes; and Rule refers to the agent’s world model, revealing the internal cognitive structures and reasoning pathways that guide decision-making or actions. Mechanism analysis posits that agent actions and interactions are the root causes of macro-level phenomena, and that mechanistic explanations clarify how these micro-level dynamics lead to emergent patterns step by step. The internal world model of an agent functions like “turning the black box into a white box”: by analyzing behavioral trajectories and cognitive chains, one can understand how individual micro-level behaviors give rise to macro-level emergent outcomes. In this context, large language models (LLMs) are powerful tools that can assist in uncovering the reasoning logic and behavioral intentions of agents.
@@ IV-A Algorithmic Mechanisms of O2O Platforms
The term “involution” was introduced by sociologist Alexander Gerschenkron to describe a state of stagnation or regression in socioeconomic development. Within social ecosystems, involution refers to excessive competition among individuals or groups in a resource-constrained environment, leading to decreased efficiency of input-output ratios and a decline in overall welfare. In this study, rider involution manifests as workers extending their working hours and increasing labor intensity in order to compete for a limited number of orders. However, their income does not increase significantly; instead, this behavior leads to deteriorating health conditions and growing income disparity. This research employs a Large Language Model (LLM) to construct a realistic multi-agent O2O platform simulation, aiming to analyze the causes and emergent dynamics of involution among delivery riders.
@@
Rules Model Rider Agent Learning and Evolution If the price per delivery order decreases, riders must work longer hours, reduce the delivery time per order, and complete more deliveries in order to maintain their income. The objective of Rider agent learning and optimization is to maximize individual utility per unit of time. Depending on the scenario, various machine learning approaches can be adopted, including large language model (LLM), evolutionary learning, and reinforcement learning (RL). On one hand, agents can utilize imitation learning to dynamically adjust the weights of their utility functions based on environmental cues and the behavior of neighboring agents. On the other hand, agents can adopt RL- or LLM-based policies during decision-making and continuously improve their strategies through interaction with the environment.
@@ IV-B Experimental Objectives and Setup
All experiments were conducted using a locally deployed DeepSeek-R1-Distill-Qwen-32B-FP8-Dynamic large language model (LLM), with the temperature parameter set to 0 to ensure repeatability. The experimental environment consists of a multi-agent system comprising five agent types: merchants, delivery riders, customers, government, and platform. For this study, we instantiated 100 rider agents.
@@ IV-C2 Intervention Analysis (RQ2)
Lastly, Figure 11(c)) demonstrates that order volume (Factor C) exerts a strong influence on the degree of involution. As the number of available orders increased, the involution index decreased steadily from 80.32 to 24.76, indicating a shift from a high-involution to a low-involution state. This finding underscores the crucial role of resource abundance—in this case, sufficient order volume—in promoting stable system operations for riders and alleviating competitive pressure within the platform.
Further analysis of the relationship between influencing factors and the degree of involution, as shown in Figure 11(d), was conducted using a structural equation model (SEM) based on linear correlation calculations. The results reveal that interaction mode transformation and order volume are the two most critical factors affecting the experimental outcomes. Their respective standardized path coefficients are 0.45 and 0.50, indicating strong and direct impacts on the observed level of involution within the system.
@@ IV-C3 Mechanism Analysis (RQ3)
To uncover the mechanisms driving these patterns, three intervention scenarios were evaluated. In the “enhanced interaction” setting (Figures 12(h)), agents overwhelmingly adopted risk-avoidant behaviors, indicating that deeper interaction intensity increases the tendency toward conservative decision-making. In the “enhanced intelligence” scenario (Figures 12(i)), the share of risk-averse intentions rose even further, suggesting that stronger cognitive capacity reinforces caution. Conversely, in the “increased order density” scenario (Figures 12(k)), the proportion of risk-aware agents dropped to 61%, implying that higher order density-induced workloads reduce risk sensitivity and lead to more risk-taking behavior.

### v2

@@ I The Laws of System Complexity
Complexity is one of the most intricate concepts in modern science [25, 39], and no unified definition has yet been established; perhaps such a definition may never exist. The different layers of the real world—physical, biological, social, and conscious layers—each exhibits distinct forms of complexity that should not be conflated. One cannot substitute lower-level complexity for higher-level complexity, nor dismiss lower-level complexity on the basis of higher-level complexity. Complexity can be understood from various perspectives and at different levels. This section first provides a categorical overview of the causes of complexity, followed by a detailed discussion using structural complexity and distributional complexity as examples. The formulas presented here are mainly drawn from Claudio Cioffi-Revilla’s work Introduction to Computational Social Science: Principles and Applications. For more details, please refer to the original text [5].
I-A Emergent Phenomena and Causal Identification Challenges in Complex Systems
@@
Complexity Arising from Individual Intelligence: Even if a system composed of non-intelligent individuals generates complexity through self-organization, it is usually relatively rudimentary and low-level. Intelligent agent systems (such as those studied by the Santa Fe Institute’s complex adaptive systems research) can sense their surrounding environments, predict future system states, learn from experience, and develop behavioral rules, thereby achieving adaptive transformations. Thus, they are inherently complex. The higher the intelligence of individual agents, the greater the system’s complexity. Agent intelligence is one of the major sources of complexity. For example, land combat systems typically involve a limited number of agents but are still complex due to the high intelligence of each agent, resulting in unpredictable and emergent system behaviors. These systems are therefore classified as open complex systems.
@@ II-A Horizontal Causation and Vertical Causation
In contrast, vertical causality argues that opening the black box involves decomposing the system into its individual components and generating the macroscopic phenomena under study by reconstructing the dynamic interactions between these components. Based on this perspective, priority should be given to methods that provide credible narratives to explain observed patterns. This view maintains that dependencies do not constitute causal mechanisms themselves, but rather are manifestations of such mechanisms. In response to this viewpoint, simulation methods become powerful tools for investigating the complex dynamic details between events. In particular, ABM is regarded as a key tool for establishing causal mechanisms, as it can demonstrate that the dependencies under study can be derived from hypothetical dynamic narratives [1].
@@
In contrast, the ”vertical causation” camp has embraced theory-driven approaches as its mainstream, requiring macroeconomic models to be integrated with ”microfoundations.” Initially, this led to the widespread adoption of rational choice models, which calculate economic aggregates based on individual preferences and expectations [19, 17]. Subsequently, ABM was introduced to address issues related to agent heterogeneity and the aggregation of social networks [42, 41, 2]. Sociology—analytical sociology in particular—has followed a similar path: shifting from regression-based statistical methods for survey data to ABM, in order to tackle micro-to-macro problems when interdependent structures exist [24].
@@
From the perspective of vertical causation, however, this implication is incorrect, because the role of variables and functions in statistical models differs from their role in ABM. In the former case, variables and functions are used to detect patterns of average effects—patterns that may reflect the overall statistical characteristics of the hypothesized underlying mechanisms (composed of entities, attributes, and interactions). In contrast, in the latter case, variables and functions are employed to precisely represent the attributes, activities, and interactions of relevant micro-level entities. Furthermore, functions are used to initiate (i.e., simulate) the dynamics associated with the hypothesized mechanisms, with the goal of assessing under what conditions such dynamics give rise to the relevant macro-level phenomena.
@@ II-B Forward Causation and Reverse Causation
Goldthorpe further clarified that causal relationships can be understood in at least three distinct ways: robust dependence, manipulability of outcomes, and generative processes [14]. In the first view, a causal claim depends on whether X continues to affect Y after introducing a set of other variables (denoted as Z) into the analysis. This perspective sees causality as contingent on controlling for confounding variables. Methods aligned with this view include time-series analysis, early causal path analysis, SEM, and more broadly, regression-based multivariate analytical techniques used in survey data analysis. In the second view, causality is interpreted as follows: if a causal factor X is manipulated, then under appropriate controls of confounding variables, a systematic effect on the response variable Y will be observed. This manipulative view of causality is aligned with the methodology of randomized experiments. Lastly, when causation is understood as a generative process, the focus shifts to explaining the nature and validity of the relational processes on which causality depends. In this view, simulation methods are considered promising tools for explaining underlying mechanisms.
@@
The third scenario belongs to the reverse causal research paradigm, characterized by a longitudinal, generative perspective, and often operationalized through ABM. In this framework, Y denotes the social phenomenon to be explained, X refers to a set of agents with diverse heterogeneous attributes, and rules specify agent behaviors and interaction mechanisms. Simulations illustrate how micro-level interactions give rise to specific macro-level phenomena. The actions and interactions of agents are the causes of emergent macro patterns. Tracing how agent-level behaviors and relationships “step-by-step” lead to these patterns provides a mechanistic explanation of the observed outcomes.
@@ III-A Computational Experiment = ABM + Experiment
In response to the challenges faced by ABM in the field of causal inference, Professor Fei-Yue Wang systematically proposed the basic ideas, concepts and methods of computational experiment method in 2004, emphasizing the circular feedback relationship between artificial systems and actual systems [44, 43, 62]. In recent years, Professor Xue and his team have carried out extensive and systematic work on computational experiment methods [55, 66, 65, 56, 64, 48, 50, 51, 52, 50, 54]. They provide a general framework for computational experiment method, which can integrate ABM logic and experimental techniques, and support the customization of general method for different application fields [47, 53, 23]. The framework of computational experiment method includes five steps: artificial society modeling, experimental system construction, computational experiment design, computational experiment analysis, and computational experiment verification, and forms a feedback closed loop [59, 58, 57, 21, 49].
Computational experiment methods are an improved and extended version of ABM. ABM is not simply used as a simulation tool, but as an “artificial social laboratory” that “spontaneously grows and cultivates” an alternative version of the actual system, and can conduct various “counterfactual experiments” on system behavior and decision analysis. On the one hand, computational experiments use ABM to emphasize vertical causality [46], pay attention to the causes and formation process of phenomena, and give causal significance to the cross-level correlation information generated by ABM (the connection between micro-mechanisms and macro-emergence). On the other hand, computational experiments “use counterfactual algorithms to construct multiple worlds” to emphasize horizontal causality, establish a probabilistic relationship between model parameters and macro-emergence through experimental mechanisms, reveal the patterns and laws within the system, and provide “reference”, “prediction” and “guidance” for the possible operation of the actual system. Because the virtual foundation on which the experimental mechanism is based, that is, the attributes, behaviors and local environment of the agent, is a replica of its real-world counterpart. Therefore, the probabilistic relationships established between different levels of analysis can be viewed as counterfactual connections across levels of analysis in the real world.
@@ III-A3 Scenario Deduction
III-A4 Mechanism-Based Explanation of Emergence
@@ III-B2 Explanatory Module: Factorial Experiment Design
III-B3 Predictive Module: Meta-Models of Artificial Society
@@ III-C3 Mechanism Analysis (Linking Micro and Macro)
The third layer represents the reverse causation research paradigm. This level involves not only analyzing the objective world model but also focusing on the agents’ subjective world models. Here, Y denotes the social phenomenon to be explained; X represents a set of agents with heterogeneous attributes; and Rule refers to the agent’s world model, revealing the internal cognitive structures and reasoning pathways that guide decision-making or actions. Mechanism analysis posits that agent actions and interactions are the root causes of macro-level phenomena, and that mechanistic explanations clarify how these micro-level dynamics lead to emergent patterns step by step. The internal world model of an agent functions like “turning the black box into a white box”: by analyzing behavioral trajectories and cognitive chains, one can understand how individual micro-level behaviors give rise to macro-level emergent outcomes. In this context, LLMs are powerful tools that can assist in uncovering the reasoning logic and behavioral intentions of agents.
@@ IV-A Algorithmic Mechanisms of O2O Platforms
The term “involution” was introduced by sociologist Alexander Gerschenkron to describe a state of stagnation or regression in socioeconomic development. Within social ecosystems, involution refers to excessive competition among individuals or groups in a resource-constrained environment, leading to decreased efficiency of input-output ratios and a decline in overall welfare. In this study, rider involution manifests as workers extending their working hours and increasing labor intensity in order to compete for a limited number of orders. However, their income does not increase significantly; instead, this behavior leads to deteriorating health conditions and growing income disparity. This research employs a LLM to construct a realistic multi-agent O2O platform simulation, aiming to analyze the causes and emergent dynamics of involution among delivery riders.
@@
Rules Model Rider Agent Learning and Evolution If the price per delivery order decreases, riders must work longer hours, reduce the delivery time per order, and complete more deliveries in order to maintain their income. The objective of Rider agent learning and optimization is to maximize individual utility per unit of time. Depending on the scenario, various machine learning approaches can be adopted, including LLM, evolutionary learning, and reinforcement learning (RL). On one hand, agents can utilize imitation learning to dynamically adjust the weights of their utility functions based on environmental cues and the behavior of neighboring agents. On the other hand, agents can adopt RL- or LLM-based policies during decision-making and continuously improve their strategies through interaction with the environment.
@@ IV-B Experimental Objectives and Setup
All experiments were conducted using a locally deployed DeepSeek-R1-Distill-Qwen-32B-FP8-Dynamic LLM, with the temperature parameter set to 0 to ensure repeatability. The experimental environment consists of a multi-agent system comprising five agent types: merchants, delivery riders, customers, government, and platform. For this study, we instantiated 100 rider agents.
@@ IV-C2 Intervention Analysis (RQ2)
Lastly, Figure 11(c) demonstrates that order volume (Factor C) exerts a strong influence on the degree of involution. As the number of available orders increased, the involution index decreased steadily from 80.32 to 24.76, indicating a shift from a high-involution to a low-involution state. This finding underscores the crucial role of resource abundance—in this case, sufficient order volume—in promoting stable system operations for riders and alleviating competitive pressure within the platform.
Further analysis of the relationship between influencing factors and the degree of involution, as shown in Figure 11(d), was conducted using a SEM based on linear correlation calculations. The results reveal that interaction mode transformation and order volume are the two most critical factors affecting the experimental outcomes. Their respective standardized path coefficients are 0.45 and 0.50, indicating strong and direct impacts on the observed level of involution within the system.
@@ IV-C3 Mechanism Analysis (RQ3)
To uncover the mechanisms driving these patterns, three intervention scenarios were evaluated. In the “enhanced interaction” setting, agents overwhelmingly adopted risk-avoidant behaviors, indicating that deeper interaction intensity increases the tendency toward conservative decision-making, as shown in Figures 12(h). In the “enhanced intelligence” scenario, the share of risk-averse intentions rose even further, suggesting that stronger cognitive capacity reinforces caution, as shown in Figures 12(i). Conversely, in the “increased order density” scenario, the proportion of risk-aware agents dropped to 61%, implying that higher order density-induced workloads reduce risk sensitivity and lead to more risk-taking behavior, as shown in Figures 12(k).